OpenStudy (anonymous):

Your parents are buying a house for $187,500. They have a good credit rating, are making a 20% down payment, and expect to pay $1,675/month. The interest rate for the mortgage is 4.85%. What must their realized income be before each month and how much interest is paid at the end of the second month? Be sure to include the following in your response: the answer to the original question and the mathematical steps for solving the problem demonstrating mathematical reasoning

OpenStudy (anonymous):

@iGreen @saifoo.khan

OpenStudy (anonymous):

First figure out how much they must borrow by getting rid of 20% of the cost of the home. Use this amount and the formula I = Prt to find how much interest was accrued in the first month. Add this amount to the borrowed total and then subtract the monthly payment. This will be the principal at the beginning of the second month, so use it in the I = Prt formula to find the interest accrued at the end of the second month. For the second part of the question, their monthly payment should be no more than 28% of their realized income, so take their payment and divide it by the 28%

OpenStudy (anonymous):

@thomaster @bohotness

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@phi @ParthKohli

OpenStudy (phi):

The good news is they gave us the instructions. The bad news is you have to understand them, and do what they say.

OpenStudy (anonymous):

ugh, thanks i been working on this question for days and it's hard to figure out.

OpenStudy (phi):

The first step is we start with $187,500 and make a "20% down payment" any idea what that means ?

OpenStudy (anonymous):

don't we have to divide that?

OpenStudy (phi):

think like this: you owe somebody $100 you pay them 20% (which means you paid them 20% of 100, and that means 20/100 * 100 ) how much did you pay ?

OpenStudy (phi):

How about you paid 1/5 of the bill i.e. you paid 1/5 of 100 how much did you pay ?

OpenStudy (anonymous):

sorry computer froze, hold on.

OpenStudy (anonymous):

so i did 20 / 100 * 100 and got 20 is that right?

OpenStudy (phi):

yes. and notice 20/100 simplifies to 1/5 so 1/5 of 100 (or 100/5 = 20, same answer)

OpenStudy (phi):

the idea is 20% is a fraction (20/100 , which may or not simplify) and you are paying some fraction of the entire bill examples: 10% of 50 is 10/100 * 50 or 1/10 * 50 = 5 how about 20% of 187500 ? any idea ?

OpenStudy (anonymous):

no. i got nothing.

OpenStudy (phi):

It's similar to the problem 1/2 of 4

OpenStudy (phi):

How would you figure out 1/2 of 4 ?

OpenStudy (anonymous):

it's 2

OpenStudy (phi):

and the way you figure it out is multiply: \[\frac{1}{2} \cdot 4 = 2 \] can you do 3/4 of 4 ?

OpenStudy (anonymous):

3

OpenStudy (phi):

now use that same idea with 20/100 of 187500 bigger numbers, but the same idea. what do you get ?

OpenStudy (anonymous):

37500

OpenStudy (phi):

that means 20% of 187500 is 37500 that is how much was paid "up front" which means was paid immediately. How much is left to pay ? any idea ?

OpenStudy (anonymous):

i have no idea

OpenStudy (phi):

if you owed 100 and paid 20, how much is left to pay ?

OpenStudy (anonymous):

80

OpenStudy (phi):

How did you figure that out ?

OpenStudy (anonymous):

well if i owed 100 and only paid 20 i subtract 100 from 20 and get 80 so i owe 80

OpenStudy (phi):

exactly. now use that same idea, but with bigger numbers you owed 187500 and paid 37500 how much is left to pay ?

OpenStudy (anonymous):

150,000

OpenStudy (phi):

that is the answer to the first question First figure out how much they must borrow by getting rid of 20% of the cost of the home. the answer is 150000

OpenStudy (anonymous):

thank you.

OpenStudy (phi):

Now the next step Use this amount and the formula I = Prt to find how much interest was accrued in the first month. when you borrow money they call the amount P (for principal) Here , P = 150000 (we just figured this out) r is the "rate" t is the "time"

OpenStudy (anonymous):

yeah p= 150000

OpenStudy (phi):

what numbers should we use for r and t ?

OpenStudy (anonymous):

i have no idea.

OpenStudy (phi):

The interest rate for the mortgage is 4.85% does that help for finding what r is ?

OpenStudy (anonymous):

yes sir.

OpenStudy (phi):

what is r ?

OpenStudy (anonymous):

so 150000 / 4.85% ?

OpenStudy (anonymous):

7275?

OpenStudy (phi):

when you see % you first change it to a fraction 4.85% becomes 4.85/100 = 0.0485 (as a decimal) next, this is rate is for a full year. we want the monthly rate so we divide by 12 (12 months in a year) that means \[ r= \frac{0.0485}{12} \] t (time) is 1 month so use t=1

OpenStudy (phi):

Here is what you know: p= 150000 r= 0.0485/12 t=1 now figure out I= p*r*t (which means multiply p times r times t, using those numbers instead of the letters)

OpenStudy (anonymous):

606.25

OpenStudy (phi):

you found I = 606.25 that is the interest that we add to the amount owed at the end of 1 month in other words, the amount owed after one month is 150000+606.25

OpenStudy (anonymous):

150606.25

OpenStudy (phi):

so far we did this: Add this amount to the borrowed total the next step is to do this: and then subtract the monthly payment.

OpenStudy (phi):

read the question to find the monthly payment what is the monthly payment ?

OpenStudy (anonymous):

148931.25

OpenStudy (phi):

yes, the payment is 1675 and we do 150606.25 - 1675= 148931.25

OpenStudy (anonymous):

yeah that is what i did.

OpenStudy (phi):

according to the question, they say ** This will be the principal at the beginning of the second month,** so after one month, and making a payment, they now owe p= 148931.25 can you figure out the interest using I = p*r*t where r and t are the same as the first time, but we have a new p ?

OpenStudy (anonymous):

601.93

OpenStudy (phi):

ok, that is the answer to this part: how much interest is paid at the end of the second month? the interest paid at the end of the second month is 601.93

OpenStudy (phi):

The last part is this: What must their realized income be? The rule is that the payment is less than or equal to 28% of their income. payment is 1675 we don't know their income, but we can give it a name: x 1675 is 28% of x or 1675= 28/100 * x to "find x" we multiply both sides by 100/28 \[ 1675 \cdot \frac{100}{28} = x \]